a chemist made 100ml of a 16% acid solution by mixing a 10% acid solution with a 30% acid solution. How much of the 30% acid solution did the chemist use
Answer by mananth(16946)

a chemist made 100ml of a 16% acid solution by mixing a 10% acid solution with a 30% acid solution. How much of the 30% acid solution did the chemist use
..
mixing a 10% acid solution : x ml
with a 30% acid solution.: 100-x ml
Mixture =100ml of a 16% acid solution
..
0.1x+0.3(100-x)=0.16*100
0.1x+30-0.3x=16
-0.2x=16-30
-0.2x=-14
/-0.2
x= -14/-0.2
x = 70 ml of 10%
...
m.ananth@hotmail.ca